OpenStudy (anonymous):

The area of a rectangle is given by A = x2 – 2x + 30. Find the value of x if the total area of the rectangle is equal to 45 square units.

OpenStudy (anonymous):

i meant x^2-2x+30

OpenStudy (cwrw238):

x^2 – 2x + 30 = 45 x^2 – 2x - 15 = 0 solve this for x and take the positive root

OpenStudy (cwrw238):

(x - 5)(x + 3) = 0 x = 5

OpenStudy (anonymous):

Thank You So Much :)

OpenStudy (cwrw238):

yw